Encyclopedia Astronautica
2008.06.07 - ISS On-Orbit Status 06/07/08


Saturday -- Flight Day 8 (FD8) of STS-124/1J.


ISS crew work cycle shifted another 30 min to the left: wake-up 6:02am EDT; sleep 9:02pm (Shuttle crew 30 min later: 9:32pm).

Crew activities aboard the ISS centered on three major areas: (1) Initial deployment of JEM RMS (Japanese Experiment Module Robotic Manipulator System) activation & checkout, (2) JLP (JEM Logistics Pressurized Module) post-relocation outfitting (Part 1), (3) Preparations for EVA-3 & EV1/EV2 Campout.

Before breakfast, FE-2 Reisman & FE-2-17 Chamitoff collected a 'wet' saliva sample (the third for Greg) for the biomed experiment INTEGRATED IMMUNE (Validating Procedures for Monitoring Crew member Immune Function). (IMMUNE protocol requires the collection to occur first thing post-sleep, before eating, drinking and brushing teeth, and all samples are stored at ambient temperature. Along with NUTRITION (Nutritional Status Assessment), INTEGRATED IMMUNE samples & analyzes participant's blood, urine, and saliva before, during and after flight for changes related to functions like bone metabolism, oxidative damage and immune function to develop and validate an immune monitoring strategy consistent with operational flight requirements and constraints. The strategy uses both long and short duration crewmembers as study subjects. The saliva is collected in two forms, dry and liquid. The dry samples are collected at intervals during the collection day using a specialized book that contains filter paper. The liquid saliva collections require that the crewmember soak a piece of cotton inside their mouth and place it in a salivette bag; there are four of the liquid collections during docked operations.)

CDR Volkov conducted the periodic service of the active Russian BMP (Harmful Impurities Removal System) by starting the "bake-out" cycle to vacuum on absorbent bed #1 of the regenerable dual-channel filtration system. The regen process will be terminated before sleeptime, at ~7:32m EDT. Regeneration of bed #2 follows tomorrow. (Regeneration of each of the two cartridges takes about 12 hours and is conducted only during crew awake periods. The BMP's regeneration cycle is regularly done every 20 days.)

Garrett Reisman, in charge of the JLP fine leak check, determined that a valve for the overnight procedure was incorrectly set. A repeat of the hermeticity check is being scheduled after Shuttle departure. (This is not a problem for ingress into JLP (on FD10), but for long term operations.)

FE-1 Kononenko conducted the periodic (currently daily) checkout/verification of IP-1 airflow sensors in the various RS (Russian Segment) hatchways, including the passageways PrK (SM Transfer Compartment) -ATV, PrK -RO (SM Working Compartment), PkhO (SM Transfer Tunnel) -RO, PkhO -DC1, PkhO -FGB PGO, FGB PGO -FGB GA, FGB GA -Node-1. (This checkup is especially important when the ventilation/circulation system has to cope with a larger crew on board, currently ten persons, and one of the two Russian SKV air conditioners off (SKV-1).)

Volkov and later Kononenko spent several hours continuing their current work with the Russian KPT-2 science payload BAR-RM, Kelvin, Ira and TTM from the RSE1 laptop, started by Oleg yesterday. (Measurements are being taken in the RS, behind panels, near welds along SM structural rings and near the shell ring in the FGB for subsequent downlinking via BSR-TM channel. At the FGB pressurized shell ring, the cosmonauts were to inspect structural elements, equipment, and cable bundles for moisture, mold, or evidence of corrosion. The data, collected at locations such as feedthroughs, windows, hatches etc., are being used to get proficient in experimenting with ISS leak detection based on environmental data anomalies (temperature, humidity, and ultrasound emissions) at potential leak locations. The BAR-RM payload uses a remote infrared thermometer (Kelvin-Video), a thermohygrometer (Iva-6A), a heat-loss anemometer/thermometer (TTM-2), an ultrasound analyzer (AU-01), and a leak detector (UT2-03) to determine physical background signs of loss of ISS pressure integrity which could be indicative of leaks in the working compartments of the station. Measurements are taken in specific zones (13 in SM PkhO and 4 in DC1), both with lights, fans & ASU pump turned on and off.)

In the FGB, Kononenko removed three IPK-1M gas masks, intended for fire suppression, and replaced them with new units. (The masks were removed along with their holding brackets on panels 230, 404 and overhead in the GA (Pressurized Adapter). The new masks were installed on Velcro in the area of the old ones.)

The FE-2 performed the regular monthly & quarterly maintenance of the TVIS (Treadmill with Vibration Isolation & Stabilization), inspecting the condition of harnesses, belt slats, corner bracket ropes, IRBAs (Isolation Restorative Bungee Assemblies) and gyroscope wire ropes for any damage or defects, lubricating as required plus recording time & date values.

For MS4 Aki Hoshide, today's activities with Kibo's own RMS (Robotic Manipulator System) began right after his wake-up and post-sleep period (~9:00am EDT), proceeding along the following steps:

Powering up RMS Monitors 1 & 2 and CCP (Camera Control Panel) by connecting to drag-thru power cable from Node-2,
Checking Joint Limit Database and powering on the Ext-2 jumper to supply electricity to RMS motor mechanisms,
Initiating (later terminating) warm-up on the #3 RMS HRM (Hold & Release Mechanism),
Release HRM,
Initial deploying of RMS, with minimal checkout of the WP (wrist pitch), then moving clear of the HRM, to allow EVA-3 spacewalkers to remove MA TVC MLIs (Main Arm/Television Camera/Multi-Layer Insulation) and LLs (Launch Locks),
Deactivating the RMS Monitors 1 & 2 and CCP for saving power resources, and
Disconnecting the drag-thru power cable.
MS1 Karen Nyberg maneuvered the SRMS (Shuttle Remote Manipulator System) to EVA-3 viewing position and later also the SSRMS (Space Station RMS) to EVA-3 egress start configuration.

Later today, JLP vestibule outfitting, Part 1, by Karen Nyberg and Aki Hoshide is involving removal of thermal covers and installation of power jumpers for JLP warm-up.

In the Lab (LAB105), Karen removed the AmiA (Antimicrobial Applicator), brought up on STS-124, from the ITCS (Internal Thermal Control System). (AmiA was used to introduce OPA (Ortho-Phthalaldehyde), an antimicrobial agent, into the Lab ITCS coolant. Later a sample of the fluid will be taken and returned for analysis.)

FE-2 Reisman filled out the regular FFQ (Food Frequency Questionnaire), his 11th, on the MEC (Medical Equipment Computer). (On the FFQs, NASA astronauts keep a personalized log of their nutritional intake over time on special MEC software. Recorded are the amounts consumed during the past week of such food items as beverages, cereals, grains, eggs, breads, snacks, sweets, fruit, beans, soup, vegetables, dairy, fish, meat, chicken, sauces & spreads, and vitamins. The FFQ is performed once a week to estimate nutrient intake from the previous week and to give recommendations to ground specialists that help maintain optimal crew health. Weekly estimation has been verified to be reliable enough that nutrients do not need to be tracked daily.)

Mark Kelly & Ken Ham again are spending several hours on cargo transfers from & to the Shuttle middeck. (As of last night, 52% of transfers were complete.)

The CDR conducted the routine daily servicing of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM, including the weekly collection of the toilet flush (SP) counter and water supply (SVO) readings for calldown to TsUP. (Regular daily SOZh maintenance consists, among else, of checking the ASU toilet facilities, replacement of the KTO & KBO solid waste containers and replacement of EDV-SV waste water and EDV-U urine containers.)

Volkov also gathered weekly data on Total Operating Time & 'On' durations of the Russian POTOK-150MK (150 micron) air filter unit of the SM's SOGS air revitalization subsystem for reporting to TsUP.

Later, Sergey completed the daily IMS (Inventory Management System) maintenance, updating/editing its standard 'delta file' including stowage locations, for the regular weekly automated export/import to its three databases on the ground (Houston, Moscow, Baikonur).

Garrett, slated for return on the Discovery on 6/14, and MS5 Greg Chamitoff, who is to replace him as ISS FE-2, had about 5 hrs scheduled for standard joint handover activities, to be continued through the docked period ahead.

The CDR conducted his third run with the Russian DZZ-2 "Diatomeya" ocean observations program, using the HDV (high-definition) video camcorder from SM windows 8 for ~30 min to record bioluminescent glow of high production zones in the South China Sea, the straits of Indonesia and the Gulf of Carpenteria (in Australia's north).

Sergey also had ~15 min set aside for another Russian "Uragan" (hurricane) earth-imaging session, using the Nikon D2X digital camera with 300-800 mm telephoto lens from SM window 9 and the RSK1 laptop for picture download. (Targets uplinked for today were the Alps, the Allaline Glacier, general views of the Carpathian mountains, forests of Central Russia Ugra National Park, water release through the Volgograd hydropower plant, and general views of the Southern Urals.)

The ISS crew completed their regular 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the CEVIS cycle ergometer (CDR, FE-2, FE-2-17), TVIS treadmill (FE-1, FE-2), RED resistive exercise device (CDR, FE-1) and VELO cycle with bungee cord load trainer (FE-1). Physical exercise regimen for Greg Chamitoff began today on the CEVIS.

Afterwards, Oleg & Greg copied the exercise data file to the MEC laptop for downlink, including the daily wristband HRM (Heart Rate Monitor) data of the workouts on RED, followed by their erasure on the HRM storage medium (done six times a week).

Fossum, Garan, Kelly, Ham, Volkov and Chamitoff conducted a one-hour review of timeline & procedures for tomorrow's EVA-3 by Mike Fossum (EV1) & Ron Garan (EV2).

Later, at ~5:30pm, after completing Airlock Equipment Lock (A/L EL) configuration for EVA-3, including recharging batteries, configuring two DCS760 cameras (one with flash for the port SARJ task) and checking EMU equipment, Mike & Ron will begin their 'campout' in the 'Quest' A/L, starting mask prebreathe, while configuring EVA tools, at ~7:57pm, then closing hatches and initiating depressurization of the CL (Crewlock) from 14.7 to 10.2 psi. Sleep for them and the ISS crew will commence at ~9:32pm, for the Shuttle crew at ~10:02pm. (The overnight Campout in the A/L CL for denitrogenation/pre-breathe at 10.2 psi lasts about 8.5 hrs. First, the two spacewalkers will perform PBA (Portable Breathing Apparatus) mask prebreathe for denitrogenation, while readying their tools & equipment, then depress the CL from 14.7 to 10.2 psi for their sleep period, to last until ~5:32am EDT tomorrow. The CL hatch will then be cracked at ~6:12am (i.e., temporarily repressurized) for a hygiene break/with mask prebreathe for Fossum & Garan. Around 7:22am, the hatch will be closed again for EVA preps in 10.2 psi, followed by EMU purge & prebreathe. Afterwards, Mark Kelly will support CL depressurization from the EL until space egress.)

EVA-3, beginning tomorrow nominally at ~10:32am EDT (likely earlier), will last an estimated 6h 30min, i.e., ending at 5:02pm. PLT Ken Ham will again be IV (Intravehicular) crewmember.
--- EVA-3 main objectives are to:

Remove old NTA (Nitrogen Tank Assembly) at S1 truss,
Retrieve spare NTA from ESP-3 (External Stowage Platform 3),
Install new NTA on S1 truss & stow old NTA,
Connect NTA FQD (Fluid Quick Disconnect) jumper,
Re-install reworked (with new power supply) ETVCG (External TV Camera Group) camera at CP9 (Camera Port 9) on the left truss,
Remove thermal covers & launch locks from JEMRMS WVE/EVE (Wrist/Elbow Vision Equipment),
Remove JPM aft window launch locks,
Deploy fwd & aft MMOD (Micrometeoroid/Orbital Debris) shields on ACBM (Active Common Berthing Mechanism), and
Collect debris samples at Port SARJ (as a get-ahead).
Waste Water Dump: At ~9:15am, PLT Ham conducted the periodic waste water dump from the Orbiter, the fourth since mission start. Two more water ventings are scheduled. (The water vent involved the waste water tank, 5 PWRs (Payload Water Reservoirs), and 2 CWCs (Contingency Water Containers). For the venting, the ISS/Shuttle stack was maneuvered from the current TEA (Torque Equilibrium Attitude) to the proper attitude (nozzles facing retrograde) and returned later to TEA.)

Port SARJ Update: Ground specialists believe the possible source of the grease-like substance observed by Fossum & Garan during EVA-2 in the portside Solar Alpha Joint is Braycote-601, a lubricant used in space, from the trundle bearings. Similar indications can be seen in imagery from the inspections performed during Mission 10A. A get-ahead task was added to tomorrow's EVA-3 to collect samples from the SARJ from debris on a tang, a second site for debris, debris mixed with grease, and the purest sample of the grease-like matter.

At ~2:07pm EDT, the STS-124 flight crew participated in three live interactive PAO TV interviews with CNN (Miles O'Brien), WCBS-TV (Erminia Vacirca), New York, NY, and WDAY-TV (Kevin Wallevand), Fargo ND.

Tonight at ~7:02pm (8:02am, 6/8, in Tokyo), the crew, Aki Hoshide & Mark Kelly, representing the crew, will receive a VIP call from Japanese Prime Minister Yasuo Fukuda, joined by Minister of Education, Culture, Sports, Science and Technology Kisaburo Tokai, Miraikan-Director and former Astronaut Dr. Mamoru Mohri, the U.S. Ambassador in Japan J. Thomas Shieffer, plus two primary school students, and two junior high school students, all located at the Miraikan (National Museum of Emerging Science and Innovation) in Tokyo.

Weekly Science Update (Expedition Seventeen -- Week 7)

ALTCRISS (Alteino Long Term monitoring of Cosmic Rays on the ISS): Measurements continue in FGB module.

ANITA (Analyzing Interferometer for Ambient Air): Continuing.

BCAT-3 (Binary Colloidal Alloy Test 3): Reserve.

CARDIOCOG-2: Completed.

CCISS (Cardiovascular & Cerebrovascular Control on Return from ISS): Reserve.

CFE (Capillary Flow Experiment): Reserve.

CW/CR (Cell Wall/Resist Wall) in EMCS (European Modular Cultivation System): In progress.

CSI-2/CGBA (CGBA Science Insert #2/Commercial Generic Bioprocessing Apparatus): In progress.

CGBA-2 (Commercial Generic Bioprocessing Apparatus 2): Complete.

CSLM-2 (Coarsening in Solid-Liquid Mixtures 2): Complete.

EarthKAM (Earth Knowledge Acquired by Middle School Students): Complete.

ELITE-S2 (Elaboratore Immagini Televisive - Space 2): Planned.

EPO (Educational Payload Operations): Reserve.

ETD (Eye Tracking Device): Completed.

EuTEF (European Technology Exposure Facility): The experiment platform continues to operate nominally. Its power feeder #1 has been de-activated to support the JEM install activities (EVA-2). The graceful shutdown of the platform was commanded some hours ahead of schedule, due to instabilities of the Ground Segment at COL-CC. After the EVA-2, the EuTEF instruments have then been re-activated nominally. A DHPU (Data Handling Processing Unit) software patch is currently developed to fix the link error issues encountered with DEBIE-2 and FIPEX instruments. - DEBIE-2: Link error still under investigation;- DOSTEL: On-going science acquisition; - EuTEMP: Currently inactive as planned; - EVC: Inactive, high rate data downlink needs further troubleshooting; - EXPOSE: On-going science acquisition; - FIPEX: A new script has been started after the re-activation of the EuTEF platform, now acquiring science; - MEDET: Acquiring science since 6/5; - PLEGPAY: Several measurements have been performed with the Langmuir probe instrument during the STS-124 (1J) docking manoeuvre. The instrument is now powered on, but not in science acquisition mode; - TRIBOLAB: The Pin On Disk run #3 (POD#3) has been nominally completed on 5/31. The instrument has been put in Thermal Stabilisation Mode after EuTEF power feeder #1 re-activation.

FSL (Fluid Science Laboratory): FSL MIL Bus cable repair was successfully performed on 5/5. After FSL Rack Activation from ground on 5/8, the ground confirmed that the FSL cables repair activities were successful. Further troubleshooting activities will be required with CEM-U (Upper) Optical Module.

GEOFLOW: Further troubleshooting activities with CEM-U(pper) Optical Module are required. A new lamp was sent up on STS-124 (1J). The start of GEOFLOW is pending further FSL troubleshooting and remaining commissioning activities (check of FSL optical modes) and it will not occur prior to the end of the 1J flight.

IMMUNO (Neuroendocrine & Immune Responses in Humans During & After Long Term Stay at ISS): Complete.

InSPACE-2 (Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions 2): In progress.

Integrated Immune: In progress.

KUBIK-FM1/ KUBIK-FM2 Centrifuge/Incubators: Completed.

LOCAD-PTS (Lab-on-a-Chip Application Development-Portable Test System): Complete.

MISSE (Materials ISS Experiment): Ongoing.

MTR-2 (Russian radiation measurements): Passive dosimeters measurements in DC1 'Pirs'.

MULTIGEN-1: Completed.

MSG-SAME (Microgravity Science Glovebox): Complete.

NOA-2 (Nitric Oxide Analyzer): Planned.

NUTRITION w/REPOSITORY: Complete.

PMDIS (Perceptual Motor Deficits in Space): Complete.

SAMS/MAMS (Space & Microgravity Acceleration Measurement Systems): Ongoing.

SAMPLE: Inc16 samples have been downloaded with 15S and handed over to science team.

SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight): 'Garrett, your only remaining activity will be to doff your Actiwatch on the Shuttle with the other two participating crewmembers. Any additional Sleep Logging on the SSC while on ISS is above and beyond and greatly appreciated by the PI. No Sleep logging is required while you are on the Shuttle.' 'Greg, you have completed your first Sleep Activity by donning the Actiwatch. You will have five download/initialization sessions and three weeks of Sleep logging remaining for your Increment 17 activities. Any additional sleep logging into the SSC will be above and beyond and greatly appreciated by the PI.'

SOLAR (Solar Monitoring Observatory): A new Sun observation period started on 6/03. Since then, the SOLAR measurements have been impacted by the ISS Attitude change and the STS-124 (1J) docking approach. Moreover, we have faced severe Telemetry and Telecommanding problems at COL-CC, leading to the postponement of science acquisition and of SOLAR criss-cross manoeuvres execution over the last days. SOLAR has been gracefully shutdown by the crew on 6/4. It was the only feasible and quick solution to safe SOVIM prior to thruster firings of the Orbiter. After the JEM activation, SOLAR power feeder #1 has been reactivated, and we are now acquiring science nominally. On 6/6, we plan to execute some criss-cross manoeuvres. - SOVIM: Acquiring science; - SOLSPEC: Acquiring science; - SOLACES: Acquiring science.

SOLO: Planned.

SPHERES (Synchronized Position Hold, Engage, Reorient, Experimental Satellite): In progress.

Swab (Characterization of Microorganisms & Allergens in Spacecraft): Complete.

3-D SPACE : A crew conference is currently planned on 6/12. First session with Greg Chamitoff is currently planned for 6/13.

TRAC (Test of Reaction & Adaptation Capabilities): Planned.

ULTRASOUND: Planned.

WAICO #1 (Waving and Coiling of Arabidopsis Roots at Different g-levels): The four ECs (Experiment Containers) of Rotor B are yet to be retrieved by the crew from the blocked Rotor B. BLB (BIOLAB) could not be activated from ground yet, pending resolution of the Smoke Detector issue. After the 1J Flight, the remaining WAICO-#1 ECs will be disposed and replaced by six Reference ECs on Rotor B.

CEO (Crew Earth Observations): Ongoing.

No CEO (Crew Earth Observation) photo targets uplinked for today.

More... - Chronology...

Home - Browse - Contact
© / Conditions for Use